Sam determines the zeros of the function f(x) to be -2 and -9.What could be Sam's function?
1. f(x)=(x-2)(x+9)
2. f(x)=(x-2)(x-9)
3. f(x)=(x+2)(x+9)
3. f(x)=(x+2)(x-9)

Found 2 solutions by math_helper, Boreal:
Answer by math_helper(2461)


Two answers are numbered 3, the top #3 (closest to #2) is correct.
If k is a root of f(x) then x-k is a factor of f(x). x-(-2) = x+2, x-(-9) = x+9

Answer by Boreal(15235)

the factors change the sign (x+2)(x+9)
The first 3.